OpenStudy (anonymous):

A motorboat can go 8 miles downstream on a river in 20 minutes. It takes 30 minutes for the boat to go upstream the same 8 miles. Find the speed of the current.

OpenStudy (anonymous):

i am prepping for my eoc exam, and still cant understand how to write theese equations, someone please help?!

OpenStudy (anonymous):

(V + s ) / 3 = 8 ( V - s) / 2 = 8